### v1

@@ Abstract
Post-quantum cryptography has gained attention due to the need for secure cryptographic systems in the face of quantum computing. Code-based and lattice-based cryptography are two prominent approaches, both heavily studied within the NIST standardization project. Code-based cryptography—most prominently exemplified by the McEliece cryptosystem—is based on the hardness of decoding random linear error-correcting codes. Despite the McEliece cryptosystem having been unbroken for several decades, it suffers from large key sizes, which has led to exploring variants using metrics than the Hamming metric, such as the Lee metric. This alternative metric may allow for smaller key sizes, but requires further analysis for potential vulnerabilities to lattice-based attack techniques. In this paper, we consider a generic Lee metric based McEliece type cryptosystem and evaluate its security against lattice-based attacks.
@@ 1 Introduction
Lattice-based cryptography, on the other hand, relies on the difficulty of solving problems in (high-dimensional) lattices, such as the Shortest Vector Problem (SVP) and the Learning With Errors (LWE) problem. Lattice-based schemes offer several compelling advantages, including strong security proofs and practical efficiency. The seminal works of Ajtai and Dwork in the late 1990s laid the groundwork for this domain, leading to the development of numerous cryptographic protocols that are both theoretically sound and practically viable.
@@ 2.1 Lee metric codes and the Lee-McEliece system
    [2, Lemma 1] Let $\mathbf{x}\in\mathbb{Z}_{q}^{n}$ be a uniformly random vector with $\mathrm{wt}_{L}\!\left(\mathbf{x}\right)=t=Tn$ for some $T\in[0,\left\lfloor q/2\right\rfloor]$. Further, let $E$ be the random variable representing a coordinate of $\mathbf{x}$. Then, as $n$ tends to infinity, for any $j\in\mathbb{Z}_{q}$,
@@
    $T=\sum_{i=0}^{q-1}\mathrm{wt}_{L}\!\left(i\right)\cdot\operatorname{Pr}(E=i).$ (2)
@@
    Let $\mathbf{x}\in\mathbb{Z}_{q}^{n}$ be a uniformly random vector with $wt_{H}(\mathbf{x})=t=\delta n$ for some $\delta\in[0,1]$. Further, let $E$ be the random variable representing a coordinate of $\mathbf{x}$. Then, as $n$ tends to infinity, for any $j\in\mathbb{Z}_{q}$,
@@
It was shown in [16] that the syndrome decoding problem is NP-complete for any additive weight function, which includes the Lee metric. It is therefore a cryptographically interesting computationally hard problem to be used in public key cryptosystems. Algorithm 1 shows a general setup of a McEliece-type public key encryption scheme with Lee metric codes.
@@ 2.2 Lattice theory
For a vector $\mathbf{r}\in\mathbb{R}^{n}$, the distance between $\mathbf{r}$ and $\mathcal{L}$ is given by $d_{1}(\mathbf{r},\mathcal{L}):=\inf\left\{d_{1}(\mathbf{r},\mathbf{v}):\mathbf{v}\in\mathcal{L}\right\}$. The shortest vector of a lattice $\mathcal{L}$ is the vector in $\mathcal{L}$ having the smallest $\ell_{1}$ norm. The length of the shortest vector is denoted by $\lambda_{1}(\mathcal{L})$, the length of the shortest lattice vector that is not a multiple of the shortest vector is denoted by $\lambda_{2}(\mathcal{L})$.
@@
[9, Theorem 1] For any $\gamma\geq 1$, there is a polynomial time reduction from $\mathsf{BDD}$1/2γ to $\mathsf{uSVP}_{\gamma}$.
@@
[9, Theorem 2] For any polynomially bounded $\gamma(n)=n^{O(1)}$, there is a polynomial time reduction from $\mathsf{uSVP}_{\gamma}$ to $\mathsf{BDD}$1/γ.
@@
The above theorem is an extension of Minkowski’s convex body theorem. Since the standard form of Minkowski’s theorem is for full-dimensional lattices and $m=1$, we provide the proof of this version in Appendix 0.A for completeness. Our proof is based on the proofs from [10, Theorem 20-21] and [14, Theorem 5-6].
@@ 2.3 Distributions
Let $F$ and $G$ be discrete probability distributions with $\text{Supp}(F)\subseteq\text{Supp}(G)$. Then,
@@
The definitions are extended in a natural way to continuous distributions using integrals instead of the summations. Note that the Kullback-Leibler divergence is a logarithm of the limit of Rényi divergence of order $a$ as $a$ goes to 1, i.e.,
@@ 3 Complexity Reductions of Lee Metric Decoding Problems
With the representation of $\mathbb{Z}_{q}^{n}$ centered around the origin, i.e., $,\mathbb{Z}_{q}^{n}=\{-\left\lfloor(q-1)/2\right\rfloor,$$\ldots,0,\ldots,\left\lfloor q/2\right\rfloor\}^{n}$, we obtain that the construction of the lattice $\mathcal{L}_{\mathsf{A}}(\mathcal{C})$ preserves the metric structure on $\mathcal{C}$, i.e., the length of the shortest $\ell_{1}$-norm vector in $\mathcal{L}_{\mathsf{A}}(\mathcal{C})$ relates to the minimum Lee distance of $\mathcal{C}$.
@@
Now, to show that $\lambda_{1}(\mathcal{L}_{\mathsf{A}})\geq\min\{q,d_{L}(\mathcal{C})\}$, it is enough to show that $\lambda_{1}(\mathcal{L}_{\mathsf{A}})\geq q$ or $\lambda_{1}(\mathcal{L}_{\mathsf{A}})\geq d_{L}(\mathcal{C})$. Let $\mathbf{x}\in\mathcal{L}_{\mathsf{A}}$ be a lattice point such that $\left\lVert\mathbf{x}\right\rVert_{1}=\lambda_{1}(\mathcal{L}_{\mathsf{A}})$. If $\mathbf{x}\bmod q={\bf 0}$, then $\lambda_{1}(\mathcal{L}_{\mathsf{A}})\geq q$ as $q$ is the smallest $\ell_{1}$-norm for a non-zero point in $\mathbb{Z}^{n}$. Now, if $\mathbf{x}\bmod q\neq{\bf 0}$, then $\mathbf{x}\in\mathcal{L}_{\mathsf{A}}\cap[-M,M]^{n}$ because, if $|x_{i}|>M$ for any $i$, then by either subtracting or adding $q$ to $x_{i}$ one can obtain another lattice point with $\ell_{1}$-norm strictly smaller than $\left\lVert\mathbf{x}\right\rVert_{1}=\lambda_{1}(\mathcal{L}_{\mathsf{A}})$, which is a contradiction. Since $\mathbf{x}\in\mathcal{L}_{\mathsf{A}}\cap[-M,M]^{n}$, we get a codeword in $\mathcal{C}$ that corresponds to $\mathbf{x}$ and has Lee weight equal to $\left\lVert\mathbf{x}\right\rVert_{1}=\lambda_{1}(\mathcal{L}_{\mathsf{A}})$. This implies that $d_{L}(\mathcal{C})\leq\lambda_{1}(\mathcal{L}_{\mathsf{A}})$.
@@
for some $\mathbf{v}\in\mathbb{Z}^{n}$ (in fact, it is $\mathbf{v}\in\{-1,0,1\}^{n}$). Then $\bar{\mathbf{c}}:=\tilde{\mathbf{c}}-\mathbf{v}q\mathbf{I}_{n}$ is an element of $\mathcal{L}_{\mathsf{A}}$ and fulfills $\tilde{\mathbf{r}}-\bar{\mathbf{c}}=\tilde{\mathbf{e}}$. This implies $d_{1}(\tilde{\mathbf{r}},\mathcal{L}_{\mathsf{A}})\leq\left\lVert\tilde{\mathbf{r}}-\bar{\mathbf{c}}\right\rVert_{1}=\left\lVert\tilde{\mathbf{e}}\right\rVert_{1}=\mathrm{wt}_{L}\!\left(\mathbf{e}\right)\leq t=\alpha\lambda_{1}(\mathcal{L}_{\mathsf{A}})$. Hence, we get an instance of $\mathsf{BDD}_{\alpha}$ for a received vector $\tilde{\mathbf{r}}$ with $d_{1}(\tilde{\mathbf{r}},\mathcal{L}_{\mathsf{A}})\leq\alpha\lambda_{1}(\mathcal{L}_{\mathsf{A}})$. The $\mathsf{BDD}_{\alpha}$ oracle now gives a lattice vector $\mathbf{x}$ satisfying $d_{1}(\tilde{\mathbf{r}},\mathbf{x})\leq\alpha\lambda_{1}(\mathcal{L}_{\mathsf{A}})=t$. Let $\mathbf{c}_{x}:=\mathbf{x}\pmod{q}$, then we have that $\mathbf{c}_{x}\in\mathcal{C}$ (according to the definition of $\mathsf{Construction\leavevmode\nobreak\ A}$ lattices) and $\mathrm{wt}_{L}\!\left(\mathbf{r}-\mathbf{c}_{x}\right)\leq t$.
@@
for some integers $c_{i,j}$. By definition, $\mathcal{L}_{\mathsf{A}}(\mathcal{C}_{\mathsf{A}}(\mathbf{B}))$ is generated by the vectors $\widetilde{\mathbf{b}_{1}},\ldots,\widetilde{\mathbf{b}_{n}},\mathbf{q}_{1},\ldots,\mathbf{q}_{n}$ so $\mathbf{b}_{i}\in\mathcal{L}_{\mathsf{A}}(\mathcal{C}_{\mathsf{A}}(\mathbf{B}))$ for all $i$, and therefore $\mathcal{L}\subseteq\mathcal{L}_{\mathsf{A}}(\mathcal{C}_{\mathsf{A}}(\mathbf{B}))$. Conversely, because $\operatorname{adj}(\mathbf{B})\cdot\mathbf{B}=q\mathbf{I}_{n}$ (or, alternatively, see [10, Theorem 16]), each $\mathbf{q}_{j}\in\mathcal{L}$ and then, for all $i$, $\widetilde{\mathbf{b}_{i}}=\mathbf{b}_{i}-\sum_{j=1}^{n}c_{i,j}\cdot\mathbf{q}_{j}\in\mathcal{L}$ so we have $\mathcal{L}_{\mathsf{A}}(\mathcal{C}_{\mathsf{A}}(\mathbf{B}))\subseteq\mathcal{L}$.
@@ 4 Containment of Finite Codes in Construction A Lattices
    if $||\mathbf{G}||_{\infty}=t\in\{0,1,\dots,M\}$ then
@@
    We can easily count the $\mathbb{Z}$-multiples of an entry $g_{i}$ in $\mathbf{G}$ within $\mathbb{Z}_{q}$ (for $\0\lambda\in\mathbb{Z}\backslash\{0\}$):
@@
    Denote by $g_{i}$ the $i$-th entry of $\mathbf{G}$. With a similar counting argument as above we get that for $\lambda\in\mathbb{Z}$, we have $|\lambda g_{i}|\leq|\lambda t|$ and
@@ 5 Comparison of Error Distributions
We can extend this distribution for length $n$ vectors over $\mathbb{Z}_{q}$ by assuming that each coordinate is independent and identically distributed and we obtain for any $\mathbf{x}\in\mathbb{Z}_{q}^{n}$:
@@
Let $F_{T}$ denote the asymptotic marginal Lee distribution and let $H_{\delta}$ denote the asymptotic marginal Hamming distribution from Lemma 2.3, both over $\mathbb{Z}_{q}$. For any $0<T<\left\lfloor q/2\right\rfloor$, let $\beta$ be the corresponding value from (2) and $c_{1}:=(\sum_{i=0}^{q-1}\exp(-\beta\mathrm{wt}_{L}\!\left(i\right)))^{-1}$. Let $\delta\in(0,1)$, then
@@
    $R_{\infty}(F_{T}||H_{\delta}\textsection)\geq c_{1}+c_{1}e^{-\beta\nu(\beta)}(q-1).$ (3)
@@
    $R_{\infty}\left(F_{T}^{(n)}||H_{\delta}^{(n)}\textsection\right)\geq\left(c_{1}+c_{1}e^{-\beta\nu(\beta)}(q-1)\right)^{n}.$
@@
Note that the first term increases as $\delta$ increases, whereas the second term decreases as $\delta$ increases. The maximum between the two minimizes when the first term is equal to the second term, i.e., $\frac{c_{1}}{1-\delta}=\frac{c_{1}e^{-\beta\nu(\beta)}(q-1)}{\delta}$ or $\delta=\frac{e^{-\beta\nu(\beta)}(q-1)}{1+e^{-\beta\nu(\beta)}(q-1)}$. Thus, we get
@@
As $q$ increases, we can observe that the distributions $F_{T}$ and $H_{\delta}$ become quite different from each other. We confirm this behavior by showing that the Rényi divergence between them goes to infinity as $q$ goes to infinity.
@@
Let $\beta$ be the corresponding value from (2) and $c_{1}:=(\sum_{i=0}^{q-1}\exp(-\beta\mathrm{wt}_{L}\!\left(i\right)))^{-1}$. As $q$ goes to infinity, we can rewrite (2) as follows:
@@
As a conclusion, we note that for a fixed $T$ (constant with respect to $q$), both $e^{-\beta}$ and $c_{1}$ converge to a constant as $q$ tends to infinity. Hence, $R_{\infty}(F_{T}||H_{\delta})\geq c_{1}+c_{1}e^{-\beta}(q-1)$ diverges as $q$ goes to infinity. This confirms the intuition that the Lee metric diverges away from the Hamming metric with growing $q$.
@@
For vectors of length $n$, we assume the distributions are independent and identical and obtain:
@@
For any given $b>0$, the KL divergence $KL(\operatorname{Lap}_{\mathbb{R},b}||\operatorname{D}_{\mathbb{R},\sigma})$ has exactly one local minimum at $\sigma=b\sqrt{2}$. In this case we obtain

### v2

@@ Abstract
Post-quantum cryptography has gained attention due to the need for secure cryptographic systems in the face of quantum computing. Code-based and lattice-based cryptography are two prominent approaches, both heavily studied within the NIST standardization project. Code-based cryptography—most prominently exemplified by the McEliece cryptosystem—is based on the hardness of decoding random linear error-correcting codes. Despite the McEliece cryptosystem having been unbroken for several decades, it suffers from large key sizes, which has led to exploring variants using metrics other than the Hamming metric, such as the Lee metric. This alternative metric may allow for smaller key sizes, but requires further analysis for potential vulnerabilities to lattice-based attack techniques. In this paper, we consider a generic Lee metric based McEliece type cryptosystem and evaluate its security against lattice-based attacks.
@@ 1 Introduction
Lattice-based cryptography, on the other hand, relies on the difficulty of solving problems in (high-dimensional) lattices, such as the Shortest Vector Problem (SVP) and the Learning With Errors (LWE) problem. Lattice-based schemes offer several compelling advantages, including strong security proofs and practical efficiency. The seminal works of Ajtai and Dwork [1] in the late 1990s laid the groundwork for this domain, leading to the development of numerous cryptographic protocols [13, 11] that are both theoretically sound and practically viable.
@@ 2.1 Lee metric codes and the Lee-McEliece system
    [3, Lemma 1] Let $\mathbf{x}\in\mathbb{Z}_{q}^{n}$ be a uniformly random vector with $\mathrm{wt}_{L}\!\left(\mathbf{x}\right)=t=Tn$ for some $T\in[0,\left\lfloor q/2\right\rfloor]$ such that $t\in\mathbb{Z}$. Further, let $E$ be the random variable representing a coordinate of $\mathbf{x}$. Then, as $n$ tends to infinity, for any $j\in\mathbb{Z}_{q}$,
@@
    $T=\sum_{j=0}^{q-1}\mathrm{wt}_{L}\!\left(j\right)\frac{\exp(-x\mathrm{wt}_{L}\!\left(j\right))}{\sum_{i=0}^{q-1}\exp(-x\mathrm{wt}_{L}\!\left(i\right))}.$ (2)
@@
    Let $\mathbf{x}\in\mathbb{Z}_{q}^{n}$ be a uniformly random vector with $wt_{H}(\mathbf{x})=t=\delta n$ for some $\delta\in[0,1]$ such that $t\in\mathbb{Z}$. Further, let $E$ be the random variable representing a coordinate of $\mathbf{x}$. Then, as $n$ tends to infinity, for any $j\in\mathbb{Z}_{q}$,
@@
It was shown in [19] that for uniformly random instances the syndrome decoding problem is NP-complete for any additive weight function, which includes the Lee metric. It is therefore a cryptographically interesting computationally hard problem to be used in public key cryptosystems. Algorithm 1 shows a general setup of a McEliece-type public key encryption scheme with Lee metric codes.
@@ 2.2 Lattice theory
For a vector $\mathbf{r}\in\mathbb{R}^{n}$, the distance between $\mathbf{r}$ and $\mathcal{L}$ is given by $d_{1}(\mathbf{r},\mathcal{L}):=\inf\left\{d_{1}(\mathbf{r},\mathbf{v}):\mathbf{v}\in\mathcal{L}\right\}$. The shortest vector of a lattice $\mathcal{L}$ is the vector in $\mathcal{L}$ having the smallest $\ell_{1}$-norm. The length of the shortest vector is denoted by $\lambda_{1}(\mathcal{L})$, the length of the shortest lattice vector that is not a multiple of the shortest vector is denoted by $\lambda_{2}(\mathcal{L})$.
@@
[10, Theorem 1] For any $\gamma\geq 1$, there is a polynomial time reduction from $\mathsf{BDD}$1/2γ to $\mathsf{uSVP}_{\gamma}$.
@@
[10, Theorem 2] For any polynomially bounded $\gamma(n)=n^{O(1)}$, there is a polynomial time reduction from $\mathsf{uSVP}_{\gamma}$ to $\mathsf{BDD}$1/γ.
@@
The above theorem is an extension of Minkowski’s convex body theorem. Since the standard form of Minkowski’s theorem is for full-dimensional lattices and $m=1$, we provide the proof of this version in Appendix 0.A for completeness. Our proof is based on the proofs from [12, Theorem 20-21] and [17, Theorem 5-6].
@@ 2.3 Distributions
Let $F$ and $G$ be discrete probability distributions satisfying $\text{Supp}(F)\subseteq\text{Supp}(G)$. Then,
@@
The definitions are extended in a natural way to continuous distributions using integrals instead of the summations. Note that we define Rényi divergence without taking the logarithm, which is standard in lattice-based cryptography. Given this, we see that the Kullback-Leibler divergence is a logarithm of the limit of Rényi divergence of order $a$ as $a$ goes to 1, i.e.,
@@ 3 Complexity Reductions of Lee Metric Decoding Problems
With the representation of $\mathbb{Z}_{q}^{n}$ centered around zero, i.e., $,\mathbb{Z}_{q}^{n}=\{-\left\lfloor(q-1)/2\right\rfloor,$$\ldots,0,\ldots,\left\lfloor q/2\right\rfloor\}^{n}$, we obtain that the construction of the lattice $\mathcal{L}_{\mathsf{A}}(\mathcal{C})$ preserves the metric structure on $\mathcal{C}$, i.e., the length of the shortest $\ell_{1}$-norm vector in $\mathcal{L}_{\mathsf{A}}(\mathcal{C})$ relates to the minimum Lee distance of $\mathcal{C}$.
@@
Now, to show that $\lambda_{1}(\mathcal{L}_{\mathsf{A}})\geq\min\{q,d_{L}(\mathcal{C})\}$, it is enough to show $\lambda_{1}(\mathcal{L}_{\mathsf{A}})\geq q$ or $\lambda_{1}(\mathcal{L}_{\mathsf{A}})\geq d_{L}(\mathcal{C})$. Let $\mathbf{x}\in\mathcal{L}_{\mathsf{A}}$ be a lattice point such that $\left\lVert\mathbf{x}\right\rVert_{1}=\lambda_{1}(\mathcal{L}_{\mathsf{A}})$. If $\mathbf{x}\bmod q={\bf 0}$, then $\lambda_{1}(\mathcal{L}_{\mathsf{A}})\geq q$ as $q$ is the smallest $\ell_{1}$-norm for a non-zero point in $\mathbb{Z}^{n}$. Now, if $\mathbf{x}\bmod q\neq{\bf 0}$, then $\mathbf{x}\in\mathcal{L}_{\mathsf{A}}\cap[-M,M]^{n}$ because, if $|x_{i}|>M$ for any $i$, then by either subtracting or adding $q$ to $x_{i}$ one can obtain another lattice point with $\ell_{1}$-norm strictly smaller than $\left\lVert\mathbf{x}\right\rVert_{1}=\lambda_{1}(\mathcal{L}_{\mathsf{A}})$, which is a contradiction. Since $\mathbf{x}\in\mathcal{L}_{\mathsf{A}}\cap[-M,M]^{n}$, we get a codeword in $\mathcal{C}$ that corresponds to $\mathbf{x}$ and has Lee weight equal to $\left\lVert\mathbf{x}\right\rVert_{1}=\lambda_{1}(\mathcal{L}_{\mathsf{A}})$. This implies that $d_{L}(\mathcal{C})\leq\lambda_{1}(\mathcal{L}_{\mathsf{A}})$.
@@
for some $\mathbf{v}\in\mathbb{Z}^{n}$ (in fact, it is $\mathbf{v}\in\{-1,0,1\}^{n}$). Here $\mathbf{I}_{n}$ denotes the identity matrix of order $n$. Then $\bar{\mathbf{c}}:=\tilde{\mathbf{c}}-\mathbf{v}q\mathbf{I}_{n}$ is an element of $\mathcal{L}_{\mathsf{A}}$ and fulfills $\tilde{\mathbf{r}}-\bar{\mathbf{c}}=\tilde{\mathbf{e}}$. This implies $d_{1}(\tilde{\mathbf{r}},\mathcal{L}_{\mathsf{A}})\leq\left\lVert\tilde{\mathbf{r}}-\bar{\mathbf{c}}\right\rVert_{1}=\left\lVert\tilde{\mathbf{e}}\right\rVert_{1}=\mathrm{wt}_{L}\!\left(\mathbf{e}\right)\leq t=\alpha\lambda_{1}(\mathcal{L}_{\mathsf{A}})$. Hence, we get an instance of $\mathsf{BDD}_{\alpha}$ for a received vector $\tilde{\mathbf{r}}$ with $d_{1}(\tilde{\mathbf{r}},\mathcal{L}_{\mathsf{A}})\leq\alpha\lambda_{1}(\mathcal{L}_{\mathsf{A}})$. The $\mathsf{BDD}_{\alpha}$ oracle now gives a lattice vector $\mathbf{x}$ satisfying $d_{1}(\tilde{\mathbf{r}},\mathbf{x})\leq\alpha\lambda_{1}(\mathcal{L}_{\mathsf{A}})=t$. Let $\mathbf{c}_{x}:=\mathbf{x}\pmod{q}$, then we have that $\mathbf{c}_{x}\in\mathcal{C}$ (according to the definition of $\mathsf{Construction\leavevmode\nobreak\ A}$ lattices) and $\mathrm{wt}_{L}\!\left(\mathbf{r}-\mathbf{c}_{x}\right)\leq t$.
@@
for some integers $c_{i,j}$. By definition, $\mathcal{L}_{\mathsf{A}}(\mathcal{C}_{\mathsf{A}}(\mathbf{B}))$ is generated by the vectors $\widetilde{\mathbf{b}_{1}},\ldots,\widetilde{\mathbf{b}_{n}},\mathbf{q}_{1},\ldots,\mathbf{q}_{n}$ so $\mathbf{b}_{i}\in\mathcal{L}_{\mathsf{A}}(\mathcal{C}_{\mathsf{A}}(\mathbf{B}))$ for all $i$, and therefore $\mathcal{L}\subseteq\mathcal{L}_{\mathsf{A}}(\mathcal{C}_{\mathsf{A}}(\mathbf{B}))$. Conversely, because $\operatorname{adj}(\mathbf{B})\cdot\mathbf{B}=q\mathbf{I}_{n}$ (or, alternatively, see [12, Theorem 16]), each $\mathbf{q}_{j}\in\mathcal{L}$ and then, for all $i$, $\widetilde{\mathbf{b}_{i}}=\mathbf{b}_{i}-\sum_{j=1}^{n}c_{i,j}\mathbf{q}_{j}\in\mathcal{L}$ so we have $\mathcal{L}_{\mathsf{A}}(\mathcal{C}_{\mathsf{A}}(\mathbf{B}))\subseteq\mathcal{L}$.
@@ 4 Containment of Finite Codes in Construction A Lattices
    if $||\mathbf{G}||_{\infty}$, the largest magnitude among the entries of $\mathbf{G}$, is equal to $t\in\{0,1,\dots,M\}$, then
@@
    Denote by $g_{i}$ the $i$-th entry of $\mathbf{G}$. We can easily count all the non-zero integer multiples of an entry $g_{i}$ that are within $\mathbb{Z}_{q}$ (for $\0\lambda\in\mathbb{Z}\backslash\{0\}$):
@@
    With a similar counting argument as above we get that for $\lambda\in\mathbb{Z}$, we have $|\lambda g_{i}|\leq|\lambda t|$ and
@@ 5 Comparison of Error Distributions
We can extend this distribution for length $n$ vectors over $\mathbb{Z}_{q}$ by assuming that each coordinate is independent and identically distributed and we obtain for any $\mathbf{x}\in\mathbb{Z}_{q}^{n}$ chosen uniformly at random:
@@
Let $F_{T}$ denote the asymptotic marginal Lee distribution and let $H_{\delta}$ denote the asymptotic marginal Hamming distribution from Lemma 2.3, both over $\mathbb{Z}_{q}$. For any $0<T<\left\lfloor q/2\right\rfloor$, let $\beta$ be the corresponding value from Equation (2) and $c_{1}:=(\sum_{i=0}^{q-1}\exp(-\beta\mathrm{wt}_{L}\!\left(i\right)))^{-1}$. Let $\delta\in(0,1)$, then
@@
    $R_{\infty}(F_{T}||H_{\delta})\geq c_{1}+c_{1}e^{-\beta\nu(\beta)}(q-1).$ (3)
@@
    $R_{\infty}\left(F_{T}^{(n)}||H_{\delta}^{(n)}\right)\geq\left(c_{1}+c_{1}e^{-\beta\nu(\beta)}(q-1)\right)^{n}.$
@@
Note that the first term increases as $\delta$ increases, whereas the second term decreases as $\delta$ increases. The maximum of the two terms minimizes when the first term is equal to the second term, i.e., $\frac{c_{1}}{1-\delta}=\frac{c_{1}e^{-\beta\nu(\beta)}(q-1)}{\delta}$ or $\delta=\frac{e^{-\beta\nu(\beta)}(q-1)}{1+e^{-\beta\nu(\beta)}(q-1)}$. Thus, we get
@@
As $q$ increases, we can observe that the distributions $F_{T}$ and $H_{\delta}$ become quite different from each other. The following result supports this observation by showing that the Rényi divergence $R_{\infty}$ between them goes to infinity as $q$ goes to infinity.
@@
From Theorem 3, we have that $R_{\infty}(F_{T}||H_{\delta})\geq c_{1}+c_{1}e^{-\beta\nu(\beta)}(q-1),$ where $c_{1}=(\sum_{i=0}^{q-1}\exp(-\beta\mathrm{wt}_{L}\!\left(i\right)))^{-1}$ and $\beta$ is the unique real solution of Equation (2), i.e.,
$T=\sum_{j=0}^{q}\mathrm{wt}_{L}\!\left(j\right)\frac{\exp(-x\ \mathrm{wt}_{L}\!\left(j\right))}{\sum_{i=0}^{q}\exp(-x\ \mathrm{wt}_{L}\!\left(i\right))}.$
Let $M=\left\lfloor q/2\right\rfloor$. Then, as $q$ goes to infinity, we can rewrite the above equation as follows:
@@
As a conclusion, we note that for a fixed $T$ (constant with respect to $q$), both $e^{-\beta}$ and $c_{1}$ converge to a constant as $q$ tends to infinity. Hence, $R_{\infty}(F_{T}||H_{\delta})\geq c_{1}+c_{1}e^{-\beta}(q-1)$ diverges as $q$ goes to infinity. This supports the intuition that the Lee metric diverges away from the Hamming metric with growing $q$.
@@
For vectors of length $n$, we assume the coordinates are independent and identically distributed, and obtain:
@@
For any given $b>0$, the KL divergence $KL(\operatorname{Lap}_{\mathbb{R},b}||\operatorname{D}_{\mathbb{R},\sigma})$ has exactly one local minimum at $\sigma=b\sqrt{2}$. In this case, we obtain